OpenStudy (anonymous):

Find ||u + v|| for u = 3 i + 3 j and v = 7 i + 3 j.

OpenStudy (anonymous):

u =<u1, u2> v =<v1,v2> (u + v ) = <u1+v1, u2+v2> =?

OpenStudy (anonymous):

be back to work, It's not hard, I don't give you the answer, period

OpenStudy (anonymous):

10,6 ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

bcoz what i did is √3^2+3^2+√7^2+3^2

OpenStudy (anonymous):

now, take magnitude of it . let say it is = v || || = sqrt((10^2) + (6^2))

OpenStudy (anonymous):

so || || =?

OpenStudy (anonymous):

2√34 ^^

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

Don't forget YOU CAN DO WHATEVER YOU LIKE. period

OpenStudy (anonymous):

im so stupid >.<

OpenStudy (anonymous):

continue your sentence: im so stupid to think that i cannot solve those easy problems. hehehe. now i change my mind, i m smart enough to pass them

OpenStudy (anonymous):

^^, thank you for your strong encouragement :)

OpenStudy (anonymous):

hihihi, i m out now. gooood luck

OpenStudy (anonymous):

can i ask last?